OpenStudy (anonymous):

the slope of a line is -4 and its y-intercept is (0,4). what is the equation of the line that's perpendicular to the first line and passes through (-8,2).

OpenStudy (e.mccormick):

Do you know the relationship between the slopes of perpendicular lines?

OpenStudy (e.mccormick):

If you do, then solve for the perpendicular slopw, then use point-slope form to solve this.

OpenStudy (anonymous):

nope. thats all the question says... our math teacher isnt that grea

OpenStudy (e.mccormick):

You should have learned the slope relationship before this point, but may not remember it. Let me do an example.

OpenStudy (anonymous):

great* and im guessing thats where i get lost casue i know the y=mx+b and the y-y1=m(x-x1) but dont know when to use them...

OpenStudy (e.mccormick):

Pkay, here are line 1 and 2, and they are perpendicular. 

OpenStudy (e.mccormick):

The slope of Line 1 is the change in y over the change in x for line 1. This is also called "Rise over Run" or "Delta Y over Delta X" In this case, the rise is 2 and the run is 3, so te slope is 2/3.

OpenStudy (e.mccormick):

With me to that point?

OpenStudy (anonymous):

yup

OpenStudy (e.mccormick):

If you look at the second line, the "rise" is not a rise at all! It is a drop of 3. For a rise to drop, it is negative. So the rise over run is -3/2. So if we look at those slopes side by side we have: \[\frac{2}{3}\, \,\mathrm{and}\;-\frac{3}{2}\] Do you see the realtionship there?

OpenStudy (e.mccormick):